### v1

# DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought

Abstract
Recently, O1-like models have emerged as representative examples, illustrating the effectiveness of long chain-of-thought (CoT) in reasoning tasks such as math and coding tasks. In this paper, we introduce DRT-o1, an attempt to bring the success of long CoT to neural machine translation (MT). Specifically, in view of the literature books that might involve similes and metaphors, translating these texts to a target language is very difficult in practice due to cultural differences. In such cases, literal translation often fails to convey the intended meaning effectively. Even for professional human translators, considerable thought must be given to preserving semantics throughout the translation process. To simulate LLMs’ long thought ability in MT, we first mine sentences containing similes or metaphors from existing literature books, and then develop a multi-agent framework to translate these sentences via long thought. In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor. To ensure the effectiveness of the long thoughts, an evaluator is also employed to judge whether the translation in the current round is better than the previous one or not. In this manner, we collect tens of thousands of long-thought MT data, which is used to train our DRT-o1. The experimental results on literature translation demonstrate the effectiveness of the DRT-o1. Using Qwen2.5-7B and Qwen2.5-14B as the backbones, the improvement brought by DRT-o1 achieves 7.33$\sim$8.26 BLEU and 1.66$\sim$3.36 CometScore. Besides, DRT-o1-7B can outperform QwQ-32B-Preview by 7.82 BLEU and 1.46 CometScore, showing its effectiveness.
DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT-o1[Uncaptioned image]DRT-o1-7B[Uncaptioned image]DRT-o1-14B
1 Introduction
Recently, the emergence of the O1-like models shows great performance in reasoning tasks, e.g., math and coding tasks OpenAI (2024b); Qin et al. (2024); Huang et al. (2024); Zhang et al. (2024); Zhao et al. (2024). With the help of long thought, LLMs tend to explore, reflect and self-improve the reasoning processes to achieve more accurate answers.
In this paper, we introduce DRT-o1 that aims to bring the success of long thought to neural machine translation (MT). There are two key points in achieving this goal:
  1. i)
    A suitable translation scenario to employ long thought in MT: Not all scenarios require long thought during translation. For example, in simple expressions, literal translation can meet most needs, and translation via long thought may be unnecessary.
  2. ii)
    A method to synthesize MT data with long thought: Long thought SFT (supervised fine-tuning) data plays a vital role in simulating LLMs’ long thought ability Huang et al. (2024). Previous work typically pays much attention to how to synthesize long-thought SFT data in math and coding tasks Qin et al. (2024); Huang et al. (2024); Zhao et al. (2024).
For i), inspired by Van den Broeck (1981), a possible scenario is translating sentences with similes or metaphors, where literal translation often fails to convey the intended semantics. Given that, we decide to mine such sentences from literature books. The mining process uses an advanced large language model (LLM) to first judge Q1: whether each literature sentence has any similes or metaphors. If has, the LLM will be asked to literally translate the sentence to a target language, and give a final judgment on Q2: whether literal translation is effective for native speakers of the target language to comprehend. If the answers of Q1 and Q2 are “yes” and “no”, respectively, the corresponding literature sentences will be reserved, and regarded as “suitable to translate via long thought”.
For ii), after collecting the literal sentences with similes or metaphors, the next question is how to synthesize long thought MT samples. Previous work typically utilizes Monte Carlo Tree Search (MCTS) Qin et al. (2024); Zhao et al. (2024); Zhang et al. (2024) or data distillation Huang et al. (2024) (from existing O1-like models) to collect long thought SFT samples. Nevertheless, MCTS is typically used in math and coding tasks where multiple reasoning behaviors should be considered, and the method emphasizes complex reasoning that might not be efficient for machine translation. Besides, utilizing existing O1-like models for data distillation might (1) constrain the potential quality of the long-thought data; and (2) have a data gap in MT since current O1-like models are typically optimized toward math and coding tasks.
Therefore, we propose a multi-agent framework to synthesize MT data with long thought. In detail, there are three agents in the framework, i.e., a translator, an advisor and an evaluator. The synthesis process is iterative, consisting of the following three steps during each iteration: (1) the translator generates a new translation conditioned on the previous step’s translation and the corresponding refinement suggestions from the advisor; (2) the advisor evaluates the current translation and offers detailed feedback; (3) the evaluator assesses the current translation and gives an evaluation score using predefined scoring criteria. Once the translation score provided by the evaluator reaches a pre-defined threshold or the number of iterations reaches a maximum value, the iteration will stop. After that, the translation and suggestions in every step could form the long-thought MT samples. To improve the readability and fluency of the long-thought data, we employ GPT-4o OpenAI (2024a) to reformulate the long-thought content.
Based on the collected long-thought MT samples, we train (SFT) our DRT-o1-7B and DRT-o1-14B using the backbones of Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct Qwen Team (2024), respectively. Experimental results on literature translation demonstrate the effectiveness of the DRT-o1. For example, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 8.26 BLEU, 1.31 CometKiwi and 3.36 CometScore. It can also outperform QwQ-32B-Preview by 7.82 BLEU and 1.46 CometScore.
Our main contributions are concluded as follows:
  • We propose DRT-o1 aiming at building LLMs with long-thought machine translation ability. To achieve this, we mine literature sentences with similes or metaphors, and collect MT samples with long-thought processes.
  • To synthesize the long-thought MT samples, we propose a multi-agent framework that involves a translator, an advisor and an evaluator. These three agents collaborate in an iterative manner to produce long thoughts during MT. Lastly, GPT-4o is used to further improve the quality of the synthetic long-thought MT samples.
  • Experimental results on literature translation verify the effectiveness of our DRT-o1. With the help of long thought, LLMs can learn to think during the machine translation.
2 DRT-o1 Data
We focus on English-to-Chinese translation, and we introduce how to collect DRT-o1 training data via three steps in this section: (1) collecting English sentences that tend to require long thoughts during translation (§ 2.1); (2) synthesizing the long-thought translation process for the collected sentences by a designed multi-agent framework (§ 2.2); (3) improving the readability and fluency of the long-thought content to form the final long-thought MT samples (§ 2.3). In the end, we provide data statistics of the collected data to give a deeper understanding (§ 2.4).
Figure 1: The illustration of the multi-agent framework to synthesize long-thought MT samples. (a) A translator iteratively produces translations under the suggestions provided by an advisor; (b) An advisor reviews the translation results and gives suggestions; (c) An evaluator assesses the translation results and gives an overall score to indicate the translation quality.
2.1 Literature Book Mining
Following Kryscinski et al. (2022), we leverage the literature books from the Project Gutenberg public-domain book repository [^1] https://www.gutenberg.org/, where the books are typically more than fifty years old and their copyrights have expired. About 400 English books are used to mine sentences with similes or metaphors.
First, we extract all sentences from these books, and filter out too short or too long sentences, i.e., less than 10 words or more than 100 words, resulting in 577.6K literature sentences. Second, for each sentence, we use Qwen2.5-72B-Instruct Qwen Team (2024) to judge whether the sentence involves similes or metaphors, and discard the sentences that do not contain any ones. Third, for the remaining sentences, we let Qwen2.5-72B-Instruct literally translate them to Chinese, and then judge whether the translation satisfies native Chinese people. If the answer is negative, the corresponding sentence will be reserved, and regarded as “suitable to translate via long thought”. In this manner, we finally collect 63K (out of 577.6K) literature sentences involving similes or metaphors whose literal translations also have flaws, called pre-collected sentences.
2.2 Multi-Agent Framework
For each pre-collected sentence (denoted as $s$), we design a multi-agent framework to translate it from English to Chinese via long thought. As shown in Figure 1, our multi-agent framework includes three agents: a translator, an advisor, and an evaluator. The synthetic process is illustrated as follows:
(1) Word-level Translation. The translator first identifies the keywords that lie in the sentence, and then provides their translations under the consideration of the context. The keywords are denoted as $\mathcal{W}^{\text{src}}=\{w^{\text{src}}_{1},w^{\text{src}}_{2},...,w^{\text{src}}_{k}\}$, where $w^{\text{src}}_{i}$ indicates the $i$-th keyword in $s$, and $k$ is the number of keywords. The translation of keywords is denoted as $\mathcal{W}^{\text{tgt}}=\{w^{\text{tgt}}_{1},w^{\text{tgt}}_{2},...,w^{\text{tgt}}_{k}\}$. This step enables the model to identify potential challenges in translating the entire sentence by breaking it down into sub-problems (i.e., word-level translation).
(2) Preliminary Translation. The translator then provides a preliminary sentence translation ($t^{0}$) conditioned on both the source sentence ($s$) and its keyword bilingual pairs ($\langle\mathcal{W}^{\text{src}},\mathcal{W}^{\text{tgt}}\rangle$).
(3) Translation Refine Loop. In the refine loop, three agents work together to refine the translation iteratively. In each iteration step $k$ (start from $k=1$), the advisor first evaluates the translation in the previous step, i.e., $t^{k-1}$, and provides detailed feedback $f^{k-1}$ for polishing it. Then, the evaluator gives an overall score of $t^{k-1}$ conditioned on both pre-defined scoring criteria and $f^{k-1}$, the score is denoted as $s^{k-1}$. In the last of the iteration step, the translator takes its previous translation $t^{k-1}$, the corresponding feedback $f^{k-1}$ and overall score $s^{k-1}$ into account to provide a new translation $t^{k}$. The translation refine loop will stop when the overall score reaches a pre-defined threshold or the number of iteration steps meets the maximum.
2.3 Long Thought Reformulation
After the multi-agent collaboration, we obtain a long thought process:
$\begin{split}\mathcal{P}(s):s\Rightarrow\langle\mathcal{W}^{\text{src}},\mathcal{W}^{\text{tgt}}\rangle\Rightarrow\langle t^{0},f^{0},s^{0}\rangle\\ \Rightarrow\langle t^{1},f^{1},s^{1}\rangle\Rightarrow...\Rightarrow\langle t^{m},f^{m},s^{m}\rangle\end{split}$ (1)
where $\mathcal{P}(s)$ denotes the multi-agent thought process for $s$, and $m$ is the number of iteration steps. To emphasize the valid thought process, translations without score change will be removed. That is, if $s^{i}$ is equal to $s^{i-1}$ ($i=1,2,...,m$), we will discard $\langle t^{i},f^{i},s^{i}\rangle$ in $\mathcal{P}(s)$, resulting in:
$\begin{split}\mathcal{P}^{{}^{\prime}}(s):s\Rightarrow\langle\mathcal{W}^{\text{src}},\mathcal{W}^{\text{tgt}}\rangle\Rightarrow\langle t^{0},f^{0},s^{0}\rangle\\ \Rightarrow\langle t^{r_{1}},f^{r_{1}},s^{r_{1}}\rangle\Rightarrow...\Rightarrow\langle t^{r_{n}},f^{r_{n}},s^{r_{n}}\rangle\end{split}$ (2)
where $1\leq r_{1}<r_{2}<...<r_{n}\leq m$, and $n$ is the number of remaining steps. If $n<3$, we will discard the whole sample, i.e., $\mathcal{P}(s)$ / $\mathcal{P}^{{}^{\prime}}(s)$.
For the remaining samples, we follow Qin et al. (2024), and leverage GPT-4o OpenAI (2024a) to modify and polish $\mathcal{P}^{{}^{\prime}}(s)$ into a self-reflection description. Finally, we obtain 22,264 machine translation samples with long thought. Figure 2 gives an example sample to illustrate the synthetic results.
Figure 2: An example of long thought synthesized by the designed multi-agent framework and GPT-4o reformulation.
# Sample Query Thought Output
o1-journey 327 41.53 486.05 3.41
Marco-O1 CoT data 10,000 52.73 673.98 52.73
DRT-o1 data ${}_{\text{(training)}}$ 19,264 37.25 527.64 44.67
DRT-o1 data ${}_{\text{(validation)}}$ 1,000 37.43 531.36 44.98
DRT-o1 data ${}_{\text{(testing)}}$ 2,000 37.19 525.44 44.70
Table 1: The number of samples and average token-level length of query, thought and output. “Query” and “Output” in DRT-o1 data mean the source sentences and the translated outputs, respectively.
2.4 Data Statistics
We split the collected 22,264 samples into training, validation and testing sets with 19,264, 1,000 and 2,000 samples, respectively. Table 1 shows the data statistics of DRT-o1 data and previous O1-like data. For Marco-O1 CoT data Zhao et al. (2024), since it is not fully released, we use its demo data to calculate the data statistics. As we can see, the average number of tokens in our synthesized thought reaches 500+ tokens, which is similar to previous math-oriented O1-like CoT data.
3 Experiments
3.1 Experimental Setups
Metrics. Following previous work, we adopt BLEU Papineni et al. (2002), reference-free CometKiwi and reference-based CometScore Rei et al. (2022) to evaluate the model translations. BLEU evaluates n-grams overlap between the generated translations and corresponding references, while CometScore evaluates the semantic similarity of translations against references. CometKiwi uses a language model to judge whether a translation conveys the semantics of the source sentence.
To calculate CometKiwi and CometScore, we leverage the official codes [^2] https://github.com/Unbabel/COMET and the official models [^3] https://huggingface.co/Unbabel/wmt22-cometkiwi-da and https://huggingface.co/Unbabel/wmt22-comet-da. To calculate the BLEU score, we use the sacrebleu toolkit [^4] https://github.com/mjpost/sacrebleu to calulate the corpus-level BLEU.
Backbones. We adopt Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct [^6] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct Yang et al. (2024) as the backbones of DRT-o1.
Figure 3: An generation case of DRT-o1-14B.
Implementation Details. Llama-Factory Zheng et al. (2024) is used to instruct-tune LLMs. Following Wang et al. (2024), all LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 32 (8$\times$4) batch size. We use the DeepSpeed optimization Rasley et al. (2020), and set ZeRO-3 optimization. Following Qin et al. (2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
Model BLEU CometKiwi CometScore
Qwen2.5-7B-Instruct 27.02 70.36 76.78
Qwen2.5-14B-Instruct 30.23 72.01 78.84
QwQ-32B-preview 27.46 71.48 78.68
Marco-o1-7B 29.48 71.62 77.41
DRT-o1-7B 35.28 71.67 80.14
DRT-o1-14B 37.56 72.16 80.50
Table 2: Experimental results on literature translation.
3.2 Main Results
Table 2 shows the results on literature translation. We compare our DRT-o1-7B and DRT-o1-14B with previous Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, QwQ-32B-preview and Marco-o1-7B. After instruction tuning on our collected data, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 8.26 BLEU, 1.31 CometKiwi and 3.36 CometScore. DRT-o1-14B outperforms Qwen2.5-14B-Instruct by 7.33 BLEU, 0.15 CometKiwi and 1.66 CometScore. Besides, DRT-o1-14B achieves the best results in terms of all metrics, showing the effectiveness of long thought in machine translation.
Figure 3 shows an example of DRT-o1-14B. As we can see, the model learns the thought process in our data collection. DRT-o1-14B first performs the word-level translation, and then attempts the preliminary translation. Next, it iteratively refines its translation until it thinks the translation is good enough.
4 Related Work
Recently, O1-like models have shown great performance in reasoning tasks, especially math and coding tasks. After the emergency of OpenAI O1 model OpenAI (2024b), many efforts are given in reproducing OpenAI O1. For example, Qin et al. (2024) propose journey learning, a training paradigm, to encourage LLMs to learn not just shortcuts, but the complete exploration process. With only 327 training samples, the journey learning enhances LLMs’ long-thought ability. Huang et al. (2024) explore the data distillation from existing O1-like models, and show the effectiveness of data distillation. Zhang et al. (2024) leverage Monte Carlo Tree Search (MCTS) to synthesize reasoning-enhanced code data, and train O1-Coder. Marco-o1 Zhao et al. (2024) is proposed to deal with open-ended text generation. Marco-o1 shows the effectiveness of a MCTS-enhanced inference method, and achieves great performance on math problems. Besides, Marco-o1 studies the machine translation cases with long thought, and shows the potentiality of O1-like models in dealing with MT.
5 Conclusion
In this paper, we introduce DRT-o1, an attempt to bring the success of long-thought reasoning to neural machine translation (MT). Specifically, we synthesize the machine translation long-thought samples by a designed multi-agent framework and GPT-4o reformulation. To collect the source sentences that are suitable for translation via long thought, we mine sentences with similes or metaphors from existing literature books. To synthesize the long thought machine translation process for these sentences, a translator, an advisor and an evaluator collaborate to translate the source sentence iteratively. GPT-4o is further employed to enhance the readability and fluency of the thought process collected via the multi-agent framework. Based on the synthesized data, we train DRT-o1-7B and DRT-o1-14B models (using Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as backbones), and show their effectiveness in literature translation.

### v2

@@
 In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor.
-To ensure the effectiveness of the long thoughts, an evaluator is also employed to judge whether the translation in the current round is better than the previous one or not.
-In this manner, we collect tens of thousands of long-thought MT data, which is used to train our DRT-o1.
-The experimental results on literature translation demonstrate the effectiveness of the DRT-o1.
-Using Qwen2.5-7B and Qwen2.5-14B as the backbones, the improvement brought by DRT-o1 achieves 7.33$\sim$8.26 BLEU and 1.66$\sim$3.36 CometScore.
-Besides, DRT-o1-7B can outperform QwQ-32B-Preview by 7.82 BLEU and 1.46 CometScore, showing its effectiveness.
+To ensure the effectiveness of the long thoughts, an evaluator is also employed to quantify the translation in each round.
+In this way, we collect tens of thousands of long-thought MT data, which is used to train our DRT-o1.
+Using Qwen2.5 and LLama-3.1 as the backbones, DRT-o1 models can learn the thought process during machine translation, and outperform vanilla LLMs as well as existing O1-like LLMs, showing their effectiveness
 DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought
-Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT-o1[Uncaptioned image]DRT-o1-7B[Uncaptioned image]DRT-o1-14B
+Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT-o1[Uncaptioned image]DRT-o1-7B[Uncaptioned image]DRT-o1-8B[Uncaptioned image]DRT-o1-14B
 1 Introduction
@@
 With the help of long thought, LLMs tend to explore, reflect and self-improve the reasoning processes to achieve more accurate answers.
-In this paper, we introduce DRT-o1 that aims to bring the success of long thought to neural machine translation (MT). There are two key points in achieving this goal:
+In this paper, our goal is not to achieve competitive performance with OpenAI’s O1 in neural machine translation (MT).
+Instead, we explore technical routes to bring the success of long thought to MT.
+To this end, we introduce DRT-o1, a byproduct of our exploration, and we hope it could facilitate the corresponding research in this direction.
+There are two key points in achieving this goal:
   1. i)
@@
 To improve the readability and fluency of the long-thought data, we employ GPT-4o OpenAI (2024a) to reformulate the long-thought content.
-Based on the collected long-thought MT samples, we train (SFT) our DRT-o1-7B and DRT-o1-14B using the backbones of Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct Qwen Team (2024), respectively.
+Based on the collected long-thought MT samples, we train (SFT) our DRT-o1-7B, DRT-o1-8B and DRT-o1-14B using the backbones of Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct Dubey et al.
+(2024) and Qwen2.5-14B-Instruct Qwen Team (2024), respectively.
 Experimental results on literature translation demonstrate the effectiveness of the DRT-o1.
-For example, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 8.26 BLEU, 1.31 CometKiwi and 3.36 CometScore.
-It can also outperform QwQ-32B-Preview by 7.82 BLEU and 1.46 CometScore.
+For example, DRT-o1-14B outperforms vanilla Qwen2.5-14B-Instruct and QwQ-32B-preview in terms of BLEU, CometKiwi, CometScore and GPT-4 evaluators.
 Our main contributions are concluded as follows:
@@
 Table 1: The number of samples and average token-level length of query, thought and output. “Query” and “Output” in DRT-o1 data mean the source sentences and the translated outputs, respectively.
+It is also worth noting that during the GPT-4o reformulation, we specify the translation with the highest score $s^{r_{j}}$ as the final translation.
+Thus, the final translation is not necessarily the last one during refinement, i.e., $t^{r_{n}}$.
+Figure 3: The distribution of the number of refinement steps in DRT-o1 data.
 2.4 Data Statistics
@@
 (2024), since it is not fully released, we use its demo data to calculate the data statistics.
-As we can see, the average number of tokens in our synthesized thought reaches 500+ tokens, which is similar to previous math-oriented O1-like CoT data.
+[^2] https://github.com/AIDC-AI/Marco-o1 As we can see, the average number of tokens in our synthesized thought reaches 500+ tokens, showing the long thought process in our data.
+Figure 3 shows the number of refinement steps/rounds in the DRT-o1 data, which ranges from three to eight.
+We can find that most samples (73.22%) involve three refinement steps, while only one sample involves eight steps.
+Furthermore, to provide a deeper understanding of the refinement process, we calculate the average edit distance before and after each refinement step.
+Specifically, the first three refinement steps cause 21.44, 13.16 and 10.90 character-level edit distance.
+This observation is consistent with intuition.
+As the refinement progresses, the magnitude of the modification gradually decreases.
+Model reference-free reference-based
+GRF CometKiwi GRB BLEU CometScore
+Vanilla LLMs
+Llama-3.1-8B-Instruct 79.25 70.14 73.30 18.55 74.58
+Qwen2.5-7B-Instruct 81.53 70.36 77.92 27.02 76.78
+Qwen2.5-14B-Instruct 84.74 72.01 80.85 30.23 78.84
+Marco-o1-7B 82.41 71.62 77.50 29.48 77.41
+QwQ-32B-preview 86.31 71.48 83.08 27.46 78.68
+DRT-o1
+DRT-o1-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 84.49 70.85 80.80 32.67 78.81
+DRT-o1-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 85.57 71.78 82.38 35.54 80.19
+DRT-o1-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 87.19 72.11 83.20 36.46 80.64
+Table 2: Experimental results on literature translation.
 3 Experiments
@@
 To calculate CometKiwi and CometScore, we leverage the official codes [^3] https://github.com/Unbabel/COMET and the official models [^4] https://huggingface.co/Unbabel/wmt22-cometkiwi-da and https://huggingface.co/Unbabel/wmt22-comet-da.
-To calculate the BLEU score, we use the sacrebleu toolkit [^4] https://github.com/mjpost/sacrebleu to calulate the corpus-level BLEU.
+To calculate the BLEU score, we use the sacrebleu toolkit [^5] https://github.com/mjpost/sacrebleu to calculate the corpus-level BLEU.
+In addition, recent studies Kocmi and Federmann (2023); Wang et al.
+(2023) also show the strong ability of LLMs in NLP evaluation.
+Thus, we use evaluators implemented using GPT-4o in reference-based and reference-free styles, which we refer to as GRB and GRF, respectively.
+Since GRB and GRF need the API costs, we randomly select 400 samples from the testing set, and conduct the GRB and GRF evaluation.
+The evaluation prompts are given in Appendix A.
 Backbones.
-We adopt Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct [^6] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct Yang et al.
-(2024) as the backbones of DRT-o1.
-Figure 3: An generation case of DRT-o1-14B.
+We adopt the following three LLMs as the backbones of our DRT-o1: (1) Llama-3.1-8B-Instruct Dubey et al.
+(2024) [^6] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct; (2) Qwen2.5-7B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and (3) Qwen2.5-14B-Instruct Yang et al.
+(2024) [^8] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct.
 Implementation Details.
@@
 Following Wang et al.
-(2024), all LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 32 (8$\times$4) batch size.
+(2024a), all LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 8 (8$\times$1) batch size.
 We use the DeepSpeed optimization Rasley et al.
@@
 (2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
-Model BLEU CometKiwi CometScore
-Qwen2.5-7B-Instruct 27.02 70.36 76.78
-Qwen2.5-14B-Instruct 30.23 72.01 78.84
-QwQ-32B-preview 27.46 71.48 78.68
-Marco-o1-7B 29.48 71.62 77.41
-DRT-o1-7B 35.28 71.67 80.14
-DRT-o1-14B 37.56 72.16 80.50
-Table 2: Experimental results on literature translation.
+Inference Details.
+When evaluating model performance on the test set, we use vLLM toolkit Kwon et al.
+(2023) [^9] https://github.com/vllm-project/vllm to accelerate the model generation.
+We use the sampling decoding strategy with 0.1 temperature, and set the repetition penalty to 1.05.
+Source Sentence Qwen2.5-14B-Instruct QwQ-32B-Preview DRT-o1-14B
+Women were weeping and children crying, and all were going as fast as seemingly lay in their power, looking behind now and then as if pursued by some deadly enemy.
+{CJK}UTF8gbsn 妇女在哭泣,孩子们在哀号,所有人都似乎在尽全力快速前行,不时回头张望,仿佛背后有某个致命的敌人在追赶。 {CJK}UTF8gbsn 妇女们哭泣着,孩子们哭喊着,每个人都在尽力加快脚步,不时回头张望,仿佛背后有致命的敌人在追击。 {CJK}UTF8gbsn 妇女悲泣孩童哀号众人疾奔不时回首,仿佛有致命之敌在后紧追。
+Wait through this season of the rain Till suns of autumn dry the plain, Then shall thy giant foe, and all His host and realm, before thee fall.
+{CJK}UTF8gbsn 等待这雨季过去,直到秋日的阳光晒干原野,那时你的巨敌及其部众和王国都将倒在你面前。 {CJK}UTF8gbsn 忍过雨季漫长夜待秋阳升空晒干大地湿润色,那时,巨敌垮台,他的军队和王国,全在你脚下 {CJK}UTF8gbsn 历尽这雨季,待秋日阳光驱散平原上的最后一丝湿润,那时,你那巨敌,连同他的军队与王国,都将匍匐在你脚下
+And if their words had taken another turn…if he himself had been less fastidious about intruding on another man’s secrets…it was cruel to think how thin a film had shut out rescue from all this guilt and misery.
+{CJK}UTF8gbsn 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 {CJK}UTF8gbsn 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 {CJK}UTF8gbsn 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
+This cold officer upon a monument, who dropped epithets unconcernedly down, would be finer as a dead man, he thought.
+{CJK}UTF8gbsn 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 {CJK}UTF8gbsn 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 {CJK}UTF8gbsn 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
+I see on your cheek two tears which I know are hot as two sparks, and salt as two crystals of the sea.
+{CJK}UTF8gbsn 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 {CJK}UTF8gbsn 我见你面颊两颗泪,温热如火花咸涩如海晶 {CJK}UTF8gbsn 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
+The night wore out, and, as he stood upon the bridge listening to the water as it splashed the river-walls of the Island of Paris, where the picturesque confusion of houses and cathedral shone bright in the light of the moon, the day came coldly, looking like a dead face out of the sky.
+{CJK}UTF8gbsn 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 {CJK}UTF8gbsn 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 {CJK}UTF8gbsn 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
+Table 3: Case Studies of literature translation. Green indicates good translations, while red indicates bad ones.
+3.2 Comparison Models
+We use vanilla Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as the comparison models.
+Besides, two O1-like LLMs, i.e., Marco-o1-7B and QwQ-32B-preview, are also conducted as baselines.
 3.3 Main Results
-Table 2 shows the results on literature translation.
-We compare our DRT-o1-7B and DRT-o1-14B with previous Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, QwQ-32B-preview and Marco-o1-7B.
-After instruction tuning on our collected data, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 8.26 BLEU, 1.31 CometKiwi and 3.36 CometScore.
-DRT-o1-14B outperforms Qwen2.5-14B-Instruct by 7.33 BLEU, 0.15 CometKiwi and 1.66 CometScore.
-Besides, DRT-o1-14B achieves the best results in terms of all metrics, showing the effectiveness of long thought in machine translation.
-Figure 3 shows an example of DRT-o1-14B.
-As we can see, the model learns the thought process in our data collection.
-DRT-o1-14B first performs the word-level translation, and then attempts the preliminary translation.
-Next, it iteratively refines its translation until it thinks the translation is good enough.
+Table 2 shows the experimental results.
+After instruction tuning on the long-thought MT training data, DRT-series LLMs significantly outperform the corresponding vanilla backbones.
+For example, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 4.04 GRF, 1.42 CometKiwi, 4.46 GRB, 8.52 BLEU and 3.41 CometScore.
+DRT-o1-8B outperforms Llama-3.1-8B-Instruct by 5.24 GRF, 0.71 CometKiwi, 7.50 GRB, 14.12 BLEU and 4.23 CometScore.
+Besides, DRT-o1-14B outperforms all baselines (including QwQ-32B-preview and Marco-o1-7B) in terms of all metrics, showing the effectiveness of long thought in MT.
+3.4 Case Study
+Table 3 provides some literature translation cases of Qwen2.5-14B-Instruct, QwQ-32B-Preview and DRT-o1-14B.
+With the help of long thought, the translations of DRT-o1-14B align more closely with the conventions of the Chinese language and exhibit a greater literary quality.
+Besides, in addition to DRT-o1-14B, some translation snippets of other LLMs can also show a great performance (marked in green).
+This indicates that vanilla LLMs might have the capability to translate literature, and long thought could further activate the capability.
+Figure 4: Time cost during inference on the testing set.
+3.5 Inference Time Analysis
+When evaluating LLMs’ literature translation performance on our test set, we leverage vLLM to accelerate the model generation.
+Specifically, we use a single NVIDIA A100 GPU (40G) to deploy and call each LLM, and calculate the corresponding generation time in the test set.
+As shown in Figure 4, the average time costs of DRT-o1 models are significantly higher than vanilla LLMs.
+This is because DRT-o1 models should first generate the long thought and then provide the final translation, thus needing more inference time.
+This also indicates that the O1-like LLMs may not be applicable to some scenarios with high real-time requirements.
 4 Related Work
-Recently, O1-like models have shown great performance in reasoning tasks, especially math and coding tasks.
+4.1 O1-like LLMs
+Recently, O1-like LLMs have shown great performance in reasoning tasks, especially math and coding tasks.
 After the emergency of OpenAI O1 model OpenAI (2024b), many efforts are given in reproducing OpenAI O1.
@@
 Besides, Marco-o1 studies the machine translation cases with long thought, and shows the potentiality of O1-like models in dealing with MT.
+4.2 Literature Translation
+Different from translating standard MT corpora (e.g., news articles), translating literature books is more difficult since it often requires equivalence beyond the word level Thai et al.
+(2022).
+Besides, it is also difficult to evaluate literature translation using automatic metrics, and previous literature translation work typically relies on human evaluation Fonteyne et al.
+(2020); Karpinska and Iyyer (2023).
+Due to its difficulty, early work is limited to small-scale attempts Genzel et al.
+(2010); Jones and Irvine (2013); Besacier and Schwartz (2015); Toral et al.
+(2018).
+Recently, Karpinska and Iyyer (2023) utilize LLMs to perform literature translation, and show that discourse-level LLM translators achieve better performances compared with sentence-level approaches.
+Thai et al.
+(2022) introduce Par3 to benchmark LLMs’ literature translation capability from non-English languages to English.
+Wang et al.
+(2024b) propose DelTA, an online document-level MT agent to translate book-level long documents based on short-term memory, long-term memory, summaries and proper noun records.
 5 Conclusion
@@
 GPT-4o is further employed to enhance the readability and fluency of the thought process collected via the multi-agent framework.
-Based on the synthesized data, we train DRT-o1-7B and DRT-o1-14B models (using Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as backbones), and show their effectiveness in literature translation.
+Based on the synthesized data, we train DRT-o1 models (using Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as backbones), and show their effectiveness in literature translation.
+Appendix A GPT-4o Evaluator
+For GRB and GRF, we prompt GPT-4o (2024-08-06 version) as the MT evaluator in the reference-based and reference-free manners, respectively.
+The corresponding prompts borrow from Kocmi and Federmann (2023), and make some adaptions to literature translation (c.f.
+Figure 5).
+Appendix B Complete Generation of DRT-o1
+Figure 6 shows an example of DRT-o1-14B.
+As we can see, the model learns the thought process in our data collection.
+DRT-o1-14B first performs the word-level translation, and then attempts the preliminary translation.
+Next, it iteratively refines its translation until it thinks the translation is good enough.

### v3

@@
-DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought
+DRT: Deep Reasoning Translation via Long Chain-of-Thought
 Abstract
 Recently, O1-like models have emerged as representative examples, illustrating the effectiveness of long chain-of-thought (CoT) in reasoning tasks such as math and coding tasks.
-In this paper, we introduce DRT-o1, an attempt to bring the success of long CoT to neural machine translation (MT).
+In this paper, we introduce DRT, an attempt to bring the success of long CoT to neural machine translation (MT).
 Specifically, in view of the literature books that might involve similes and metaphors, translating these texts to a target language is very difficult in practice due to cultural differences.
@@
 In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor.
-To ensure the effectiveness of the long thoughts, an evaluator is also employed to quantify the translation in each round.
-In this way, we collect tens of thousands of long-thought MT data, which is used to train our DRT-o1.
-Using Qwen2.5 and LLama-3.1 as the backbones, DRT-o1 models can learn the thought process during machine translation, and outperform vanilla LLMs as well as existing O1-like LLMs, showing their effectiveness
-DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought
-Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT-o1[Uncaptioned image]DRT-o1-7B[Uncaptioned image]DRT-o1-8B[Uncaptioned image]DRT-o1-14B
+To ensure the effectiveness of the long thoughts, an evaluator is also employed to quantify the translation quality in each round.
+In this way, we collect tens of thousands of long-thought MT data, which is used to train our DRT.
+Using Qwen2.5 and LLama-3.1 as the backbones, DRT models can learn the thought process during machine translation, and outperform vanilla LLMs as well as LLMs which are simply fine-tuning on the paired sentences without long thought, showing its effectiveness.
+DRT: Deep Reasoning Translation via Long Chain-of-Thought
+Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT[Uncaptioned image]DRT-7B[Uncaptioned image]DRT-8B[Uncaptioned image]DRT-14B
 1 Introduction
-Recently, the emergence of the O1-like models shows great performance in reasoning tasks, e.g., math and coding tasks OpenAI (2024b); Qin et al.
+Recently, the emergence of the O1-like LLMs shows great performance in reasoning tasks, e.g., math and coding tasks OpenAI (2024b); Qin et al.
 (2024); Huang et al.
@@
 With the help of long thought, LLMs tend to explore, reflect and self-improve the reasoning processes to achieve more accurate answers.
-In this paper, our goal is not to achieve competitive performance with OpenAI’s O1 in neural machine translation (MT).
-Instead, we explore technical routes to bring the success of long thought to MT.
-To this end, we introduce DRT-o1, a byproduct of our exploration, and we hope it could facilitate the corresponding research in this direction.
+In this paper, we explore technical routes to bring the success of long thought to MT.
+To this end, we introduce DRT, a product of our exploration, and we hope it could facilitate the research community.
 There are two key points in achieving this goal:
-  1. i)
-    A suitable translation scenario to employ long thought in MT: Not all scenarios require long thought during translation.
-For example, in simple expressions, literal translation can meet most needs, and translation via long thought may be unnecessary.
-  2. ii)
-    A method to synthesize MT data with long thought: Long thought SFT (supervised fine-tuning) data plays a vital role in simulating LLMs’ long thought ability Huang et al.
+i) A suitable translation scenario to employ long thought in MT: Not all scenarios require long chain-of-thought (CoT)111“long CoT” is equal to “long thought”, and we alternatively use these two terms in this paper.
+during translation.
+For example, in simple expressions, literal translation can meet most needs, and translation via long CoT may be unnecessary.
+Inappropriate scenarios might cause the overthinking issue Chen et al.
 (2024).
-Previous work typically pays much attention to how to synthesize long-thought SFT data in math and coding tasks Qin et al.
+ii) A method to synthesize MT data with long thought: Long thought SFT (supervised fine-tuning) data plays a vital role in simulating LLMs’ long thought ability Huang et al.
+(2024).
+Previous work pays much attention to how to synthesize long-thought data in math and coding tasks Qin et al.
 (2024); Huang et al.
@@
 To improve the readability and fluency of the long-thought data, we employ GPT-4o OpenAI (2024a) to reformulate the long-thought content.
-Based on the collected long-thought MT samples, we train (SFT) our DRT-o1-7B, DRT-o1-8B and DRT-o1-14B using the backbones of Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct Dubey et al.
-(2024) and Qwen2.5-14B-Instruct Qwen Team (2024), respectively.
-Experimental results on literature translation demonstrate the effectiveness of the DRT-o1.
-For example, DRT-o1-14B outperforms vanilla Qwen2.5-14B-Instruct and QwQ-32B-preview in terms of BLEU, CometKiwi, CometScore and GPT-4 evaluators.
+Based on the collected long-thought MT samples, we train our DRT-7B, DRT-8B and DRT-14B using the backbones of Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct Dubey et al.
+(2024) and Qwen2.5-14B-Instruct Yang et al.
+(2024a), respectively.
+Experimental results on literature translation verify their effectiveness.
+In particular, DRT-14B outperforms QwQ-32B-preview and DeepSeek-R1-Distill-Qwen-32B in terms of BLEU, CometKiwi, CometScore and GPT-4 evaluations.
+Moreover, human evaluation and case study show the strong translation performance of DRT models.
 Our main contributions are concluded as follows:
-  • We propose DRT-o1 aiming at building LLMs with long-thought machine translation ability.
+  • We propose DRT aiming at building LLMs with long-thought machine translation ability.
 To achieve this, we mine literature sentences with similes or metaphors, and collect MT samples with long-thought processes.
@@
 These three agents collaborate in an iterative manner to produce long thoughts during MT.
-Lastly, GPT-4o is used to further improve the quality of the synthetic long-thought MT samples.
-  • Experimental results on literature translation verify the effectiveness of our DRT-o1. With the help of long thought, LLMs can learn to think during the machine translation.
-2 DRT-o1 Data
-We focus on English-to-Chinese translation, and we introduce how to collect DRT-o1 training data via three steps in this section: (1) collecting English sentences that tend to require long thoughts during translation (§ 2.1); (2) synthesizing the long-thought translation process for the collected sentences by a designed multi-agent framework (§ 2.2); (3) improving the readability and fluency of the long-thought content to form the final long-thought MT samples (§ 2.3).
-In the end, we provide data statistics of the collected data to give a deeper understanding (§ 2.4).
+Lastly, GPT-4o is used to further improve the quality of the synthesized long-thought MT samples.
+  • Experimental results on literature translation verify the effectiveness of our DRT. With the help of long thought, LLMs can learn to think during the machine translation.
+2 DRT Data
+We focus on English-to-Chinese translation [^2] Although we focus on English-to-Chinese translation in this work, the methods we introduced can be trivially applied to other languages or translation directions., and we introduce how to collect the long-thought MT samples via three steps in this section: (1) collecting English sentences that tend to require long thoughts during translation (§ 2.1); (2) synthesizing the long-thought translation process for the collected sentences by a designed multi-agent framework (§ 2.2); (3) improving the readability and fluency of the long-thought content to form the final long-thought MT samples (§ 2.3).
+Next, we provide data statistics and data analyses of the collected data to give a deeper understanding (§ 2.4).
+Finally, we discuss the data quality (§ 2.5).
 Figure 1: The illustration of the multi-agent framework to synthesize long-thought MT samples.
@@
 First, we extract all sentences from these books, and filter out too short or too long sentences, i.e., less than 10 words or more than 100 words, resulting in 577.6K literature sentences.
-Second, for each sentence, we use Qwen2.5-72B-Instruct Qwen Team (2024) to judge whether the sentence involves similes or metaphors, and discard the sentences that do not contain any ones.
+Second, for each sentence, we use Qwen2.5-72B-Instruct Yang et al.
+(2024a) to judge whether the sentence involves similes or metaphors, and discard the sentences that do not contain any ones.
 Third, for the remaining sentences, we let Qwen2.5-72B-Instruct literally translate them to Chinese, and then judge whether the translation satisfies native Chinese people.
 If the answer is negative, the corresponding sentence will be reserved, and regarded as “suitable to translate via long thought”.
-In this manner, we finally collect 63K (out of 577.6K) literature sentences involving similes or metaphors whose literal translations also have flaws, called pre-collected sentences.
+For prompt details, please refer to Appendix A.1.
+Consequently, we collect 63K (out of 577.6K) literature sentences involving similes or metaphors whose literal translations have flaws, called pre-collected sentences.
 2.2 Multi-Agent Framework
-For each pre-collected sentence (denoted as $s$), we design a multi-agent framework to translate it from English to Chinese via long thought.
-As shown in Figure 1, our multi-agent framework includes three agents: a translator, an advisor, and an evaluator.
+For each pre-collected sentence (denoted as $s$), we design a multi-agent framework to translate it via long thought.
+As shown in Figure 1, our multi-agent framework includes three agents: a translator, an advisor, and an evaluator, each of which use Qwen2.5-72B-Instruct as the backbone.
 The synthetic process is illustrated as follows:
@@
 In each iteration step $k$ (start from $k=1$), the advisor first evaluates the translation in the previous step, i.e., $t^{k-1}$, and provides detailed feedback $f^{k-1}$ for polishing it.
-Then, the evaluator gives an overall score of $t^{k-1}$ conditioned on both pre-defined scoring criteria and $f^{k-1}$, the score is denoted as $s^{k-1}$.
+Then, the evaluator gives an overall score of $t^{k-1}$ conditioned on both pre-defined scoring criteria and $f^{k-1}$, and the score is denoted as $s^{k-1}$.
 In the last of the iteration step, the translator takes its previous translation $t^{k-1}$, the corresponding feedback $f^{k-1}$ and overall score $s^{k-1}$ into account to provide a new translation $t^{k}$.
 The translation refine loop will stop when the overall score reaches a pre-defined threshold or the number of iteration steps meets the maximum.
+For prompt details of the translator, advisor and evaluator, please refer to Appendix A.2.
+Figure 2: An example of long thought synthesized by the designed multi-agent framework and GPT-4o reformulation.
 2.3 Long Thought Reformulation
@@
 $\begin{split}\mathcal{P}^{{}^{\prime}}(s):s\Rightarrow\langle\mathcal{W}^{\text{src}},\mathcal{W}^{\text{tgt}}\rangle\Rightarrow\langle t^{0},f^{0},s^{0}\rangle\\ \Rightarrow\langle t^{r_{1}},f^{r_{1}},s^{r_{1}}\rangle\Rightarrow...\Rightarrow\langle t^{r_{n}},f^{r_{n}},s^{r_{n}}\rangle\end{split}$ (2)
-where $1\leq r_{1}<r_{2}<...<r_{n}\leq m$, and $n$ is the number of remaining steps. If $n<3$, we will discard the whole sample, i.e., $\mathcal{P}(s)$ / $\mathcal{P}^{{}^{\prime}}(s)$.
+where $1\leq r_{1}<r_{2}<...<r_{n}\leq m$, and $n$ is the number of remaining steps. If $n<3$, we will discard the whole sample, i.e., $\mathcal{P}(s)$.
 For the remaining samples, we follow Qin et al.
-(2024), and leverage GPT-4o OpenAI (2024a) to modify and polish $\mathcal{P}^{{}^{\prime}}(s)$ into a self-reflection description.
-Finally, we obtain 22,264 machine translation samples with long thought.
+(2024), and leverage GPT-4o OpenAI (2024a) to modify and polish $\mathcal{P}^{{}^{\prime}}(s)$ into a self-reflection description (the used prompt is provided in Appendix A.3).
+Finally, we obtain 22,264 MT samples with long thought.
 Figure 2 gives an example sample to illustrate the synthetic results.
-Figure 2: An example of long thought synthesized by the designed multi-agent framework and GPT-4o reformulation.
 # Sample Query Thought Output
@@
 Marco-O1 CoT data 10,000 52.73 673.98 52.73
-DRT-o1 data ${}_{\text{(training)}}$ 19,264 37.25 527.64 44.67
-DRT-o1 data ${}_{\text{(validation)}}$ 1,000 37.43 531.36 44.98
-DRT-o1 data ${}_{\text{(testing)}}$ 2,000 37.19 525.44 44.70
-Table 1: The number of samples and average token-level length of query, thought and output. “Query” and “Output” in DRT-o1 data mean the source sentences and the translated outputs, respectively.
+DRT data ${}_{\text{(training)}}$ 19,264 37.25 527.64 44.67
+DRT data ${}_{\text{(validation)}}$ 1,000 37.43 531.36 44.98
+DRT data ${}_{\text{(testing)}}$ 2,000 37.19 525.44 44.70
+Table 1: The number of samples and average token-level length of query, thought and output. “Query” and “Output” in DRT data mean the source sentences and the translated outputs, respectively.
 It is also worth noting that during the GPT-4o reformulation, we specify the translation with the highest score $s^{r_{j}}$ as the final translation.
 Thus, the final translation is not necessarily the last one during refinement, i.e., $t^{r_{n}}$.
-Figure 3: The distribution of the number of refinement steps in DRT-o1 data.
-2.4 Data Statistics
+Figure 3: The distribution of the number of refinement steps in DRT data.
+2.4 Data Statistics and Data Analyses
+Figure 4: Trends in average scores (provided by the evaluator agent) over the refinement steps.
+The trends for samples with three, four, five, and six refinement steps are illustrated in (a), (b), (c), and (d), respectively.
 We split the collected 22,264 samples into training, validation and testing sets with 19,264, 1,000 and 2,000 samples, respectively.
-Table 1 shows the data statistics of DRT-o1 data and previous O1-like data.
+Table 1 shows the data statistics of DRT data and previous O1-like data.
 For Marco-O1 CoT data Zhao et al.
@@
 [^4] https://github.com/AIDC-AI/Marco-o1 As we can see, the average number of tokens in our synthesized thought reaches 500+ tokens, showing the long thought process in our data.
-Figure 3 shows the number of refinement steps/rounds in the DRT-o1 data, which ranges from three to eight.
-We can find that most samples (73.22%) involve three refinement steps, while only one sample involves eight steps.
+Refine Loop Analyses.
+Figure 3 shows the number of refinement steps in the DRT data, which ranges from 3 to 8 steps.
+We can find that most samples (73.22%) involve 3 refinement steps, while only one sample involves 8 steps.
 Furthermore, to provide a deeper understanding of the refinement process, we calculate the average edit distance before and after each refinement step.
@@
 As the refinement progresses, the magnitude of the modification gradually decreases.
+To further understand the improvement brought by the translation refine loop, we calculate the average overall scores (provided by the evaluator agent) along with each refinement step.
+As shown in Figure 4, as the number of refinement steps increases, the average score generally increases, demonstrating that the refine loop could iteratively increase the quality of translations.
+Metric ACC. (%)
+CometKiwi 56.0
+Evaluator Agent (Qwen2.5-72B-Instruct) 92.5
+Evaluator Agent (GPT-4o) 93.5
+Table 2: Accuracy of automatic metrics for translation quality estimation (ACC.: accuracy).
+2.5 Quality Analyses
+The Effectiveness of the Evaluator Agent.
+Previous work has shown that the state-of-the-art LLMs can be used as evaluators for various text generation tasks Kocmi and Federmann (2023); Wang et al.
+(2023); Li et al.
+(2024).
+To figure out the effectiveness of our evaluator agent, we randomly select 200 source sentences from DRT data, and for each of them, we further select its two translations as well as scores (provided by the evaluator agent) during refinement.
+We next employ human annotators to compare the two translations of each source sentence, and judge which translation is better, or two translations are similar in quality (annotation details can be found in Appendix B).
+After obtaining the quality labels, we calculate the accuracy of the evaluator agent according to its evaluation score.
+For comparison, we also calculate the accuracy of CometKiwi Rei et al.
+(2022) and GPT-4o evaluator agent.
+As shown in Table 2, our evaluator agent achieves a high accuracy (92.5%), demonstrating its effectiveness in evaluating literature translation quality.
+Besides, the widely-used CometKiwi metric only achieves 56.0% accuracy.
+Thought CometKiwi is powerful in the general domain (e.g., news) Kocmi and Federmann (2023), its effectiveness in the literature domain is limited and unreliable, which is also pointed out by Karpinska and Iyyer (2023).
+Furthermore, the GPT-4o evaluator agent slightly outperforms the origin evaluator agent (with Qwen2.5-72B-Instruct backbone).
+Considering the tradeoff between cost and effectiveness, we finally decide to use Qwen2.5-72B-Instruct as our evaluator agent.
+Translation Quality.
+Based on the effectiveness of the evaluator agent and the observation that evaluation scores of final translations typically reach 90.0 (c.f., Figure 4), we can ensure a high level of translation quality in the constructed data.
+According to the pre-defined scoring criteria of the evaluator agent (c.f., Appendix A.2), a score of 90.0 indicates excellent translations.
 Model reference-free reference-based
-GRF CometKiwi GRB BLEU CometScore
+GEA GRF CometKiwi GRB BLEU CometScore
 Vanilla LLMs
-Llama-3.1-8B-Instruct 79.25 70.14 73.30 18.55 74.58
-Qwen2.5-7B-Instruct 81.53 70.36 77.92 27.02 76.78
-Qwen2.5-14B-Instruct 84.74 72.01 80.85 30.23 78.84
-Marco-o1-7B 82.41 71.62 77.50 29.48 77.41
-QwQ-32B-preview 86.31 71.48 83.08 27.46 78.68
-DRT-o1
-DRT-o1-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 84.49 70.85 80.80 32.67 78.81
-DRT-o1-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 85.57 71.78 82.38 35.54 80.19
-DRT-o1-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 87.19 72.11 83.20 36.46 80.64
+Llama-3.1-8B-Instruct 59.58 79.25 70.14 73.30 18.55 74.58
+Qwen2.5-7B-Instruct 66.21 81.53 70.36 77.92 27.02 76.78
+Qwen2.5-14B-Instruct 70.86 84.74 72.01 80.85 30.23 78.84
+Marco-o1-7B 64.24 82.41 71.62 77.50 29.48 77.41
+QwQ-32B-preview 75.50 86.31 71.48 83.08 27.46 78.68
+DeepSeek-R1-Distill-Llama-8B 56.89 76.31 67.13 69.49 15.83 71.82
+DeepSeek-R1-Distill-Qwen-7B 43.66 65.16 63.49 58.13 10.99 69.21
+DeepSeek-R1-Distill-Qwen-14B 70.64 83.92 71.01 80.29 25.55 77.66
+DeepSeek-R1-Distill-Qwen-32B 71.88 84.78 71.93 81.59 29.36 78.93
+SFT LLMs (w/o CoT)
+Llama-3.1-8B-SFT 69.33 84.10 70.25 80.18 30.03 78.26
+Qwen2.5-7B-SFT 72.29 85.06 71.03 81.72 35.44 80.10
+Qwen2.5-14B-SFT 74.53 85.66 72.08 83.08 37.63 80.82
+DRT
+DRT-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
+DRT-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
+DRT-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
 Table 3: Experimental results on literature translation.
+The bold and the underline denote the best and second-best performances, respectively.
+$\dagger$” and“$\ddagger$” denote statistically significant better than the corresponding SFT LLMs (w/o CoT) with t-test p < 0.01 and 0.05, respectively.
 3 Experiments
@@
 Metrics.
-Following previous work, we adopt BLEU Papineni et al.
-(2002), reference-free CometKiwi and reference-based CometScore Rei et al.
+Following previous work, we adopt “BLEU” Papineni et al.
+(2002), “CometKiwi” and “CometScore” Rei et al.
 (2022) to evaluate the model translations.
-BLEU evaluates n-grams overlap between the generated translations and corresponding references, while CometScore evaluates the semantic similarity of translations against references.
-CometKiwi uses a language model to judge whether a translation conveys the semantics of the source sentence.
-To calculate CometKiwi and CometScore, we leverage the official codes [^3] https://github.com/Unbabel/COMET and the official models [^4] https://huggingface.co/Unbabel/wmt22-cometkiwi-da and https://huggingface.co/Unbabel/wmt22-comet-da.
-To calculate the BLEU score, we use the sacrebleu toolkit [^5] https://github.com/mjpost/sacrebleu to calculate the corpus-level BLEU.
-In addition, recent studies Kocmi and Federmann (2023); Wang et al.
-(2023) also show the strong ability of LLMs in NLP evaluation.
-Thus, we use evaluators implemented using GPT-4o in reference-based and reference-free styles, which we refer to as GRB and GRF, respectively.
-Since GRB and GRF need the API costs, we randomly select 400 samples from the testing set, and conduct the GRB and GRF evaluation.
-The evaluation prompts are given in Appendix A.
+Among them, BLEU evaluates n-grams overlap between model translations and references, while CometScore evaluates the semantic similarity of model translations against references.
+CometKiwi uses a language model to judge whether a model translation conveys the semantics of the source sentence.
+As pointed out by Karpinska and Iyyer (2023), BLEU and COMET may be ineffective for evaluating literature translation.
+Meanwhile, recent studies also show the strong ability of LLMs in NLP evaluation Li et al.
+(2024).
+Therefore, we use evaluators implemented using GPT-4o in reference-based and reference-free styles, which we refer to as “GRB” and “GRF”, respectively.
+The evaluation prompts borrow from Kocmi and Federmann (2023), and are illustrated in Appendix C.
+Furthermore, as demonstrated in § 2.4, the GPT-4o evaluator agent achieves great accuracy in literature translation.
+We also leverage it as the evaluation metric in experiments, which is referred to as “GEA”.
+Since GRB, GRF and GEA need the API costs, we randomly select 400 samples to conduct evaluation.
 Backbones.
-We adopt the following three LLMs as the backbones of our DRT-o1: (1) Llama-3.1-8B-Instruct Dubey et al.
-(2024) [^6] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct; (2) Qwen2.5-7B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and (3) Qwen2.5-14B-Instruct Yang et al.
-(2024) [^8] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct.
-Implementation Details.
-Llama-Factory Zheng et al.
-(2024) is used to instruct-tune LLMs.
-Following Wang et al.
-(2024a), all LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 8 (8$\times$1) batch size.
-We use the DeepSpeed optimization Rasley et al.
-(2020), and set ZeRO-3 optimization.
-Following Qin et al.
-(2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
-Inference Details.
-When evaluating model performance on the test set, we use vLLM toolkit Kwon et al.
-(2023) [^9] https://github.com/vllm-project/vllm to accelerate the model generation.
-We use the sampling decoding strategy with 0.1 temperature, and set the repetition penalty to 1.05.
-Source Sentence Qwen2.5-14B-Instruct QwQ-32B-Preview DRT-o1-14B
-Women were weeping and children crying, and all were going as fast as seemingly lay in their power, looking behind now and then as if pursued by some deadly enemy.
-{CJK}UTF8gbsn 妇女在哭泣,孩子们在哀号,所有人都似乎在尽全力快速前行,不时回头张望,仿佛背后有某个致命的敌人在追赶。 {CJK}UTF8gbsn 妇女们哭泣着,孩子们哭喊着,每个人都在尽力加快脚步,不时回头张望,仿佛背后有致命的敌人在追击。 {CJK}UTF8gbsn 妇女悲泣孩童哀号众人疾奔不时回首,仿佛有致命之敌在后紧追。
-Wait through this season of the rain Till suns of autumn dry the plain, Then shall thy giant foe, and all His host and realm, before thee fall.
-{CJK}UTF8gbsn 等待这雨季过去,直到秋日的阳光晒干原野,那时你的巨敌及其部众和王国都将倒在你面前。 {CJK}UTF8gbsn 忍过雨季漫长夜待秋阳升空晒干大地湿润色,那时,巨敌垮台,他的军队和王国,全在你脚下 {CJK}UTF8gbsn 历尽这雨季,待秋日阳光驱散平原上的最后一丝湿润,那时,你那巨敌,连同他的军队与王国,都将匍匐在你脚下
+We adopt the following three LLMs as the backbones of our DRT: Llama-3.1-8B-Instruct Dubey et al.
+(2024), Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct Yang et al.
+(2024b).
+All model checkpoints are publicly available.
+For evaluation toolkits and the implementation details of all models, please refer to Appendix D.
+3.2 Comparison Models
+Vanilla LLMs.
+We leverage vanilla Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct Yang et al.
+(2024a) as the comparison models.
+Besides, six O1-like LLMs are also conducted as baselines: Marco-o1-7B Zhao et al.
+(2024), QwQ-32B-preview Qwen (2024), DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-14B and DeepSeek-R1-Distill-Qwen-32B Guo et al.
+(2025).
+SFT LLMs (w/o CoT).
+We also fine-tune LLMs with only paired sentences of DRT training data (without thought).
+This setting allows LLMs to learn the mapping from source literature sentences to the corresponding Chinese translations directly.
+We denote the fine-tuned LLMs as Llama-3.1-8B-SFT, Qwen2.5-7B-SFT and Qwen2.5-14B-SFT, serving as strong baselines in the experiments.
+3.3 Main Results
+Table 3 shows the experimental results, we analyze the performance from the following aspects:
+SFT LLMs (w/o CoT) vs.
+Vanilla LLMs.
+After instruction tuning on the paired sentences of our training data, SFT LLMs (w/o CoT) significantly outperform the corresponding vanilla LLMs.
+For example, Llama-3.1-8B-SFT outperforms Llama-3.1-8B-Instruct by 9.75 GEA, 4.85 GRF and 6.88 GRB.
+Qwen2.5-7B-SFT outperforms Qwen2.5-7B-Instruct by 6.08 GEA, 3.53 GRF and 3.80 GRB.
+This finding demonstrates the effectiveness of our multi-agent framework and the quality of the synthesized translation.
+Please also note that the final translations are synthesized by Qwen2.5-72B-Instruct, indicating that we can leverage off-the-shelf open-source LLMs to collect high-quality literation translation data.
+And the data could help smaller LLMs (such as 7B and 14B ones) to boost their literature translation skills.
+DRT vs.
+Vanilla LLMs.
+After fine-tuning on the long-thought MT training data, our DRT-series LLMs also significantly outperform the corresponding vanilla backbones.
+Particularly, DRT-14B outperforms QwQ-32B-preview and DeepSeek-R1-Distill-Qwen-32B in terms of all metrics, showing its effectiveness in literature MT.
+DRT vs.
+SFT LLMs (w/o CoT).
+Using Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct as backbones, LLMs tuned with long thought achieve better performance than those tuned without long thought in terms of all metrics.
+For example, DRT-7B outperforms Qwen2.5-7B-SFT by 2.76 GEA, 0.51 GRF, 0.75 CometKiwi, 0.66 GRB, 0.10 BLEU and 0.09 CometScore.
+When using Qwen2.5-14B-Instruct as the backbone, we find that DRT-14B outperforms Qwen2.5-14B-SFT in terms of GEA, GRF, CometKiwi and GRB, but underperforms in terms of BLEU and CometScore.
+In detail, BLEU and CometScore evaluate the translations from the perspective of similarity between model translations and golden references.
+We conjecture that the higher BLEU and CometScore performance of Qwen2.5-14B-SFT is due to the model’s ability to quickly learn domain-specific translations through tuning without long thoughts, allowing it to adapt to the literature translation more straightforwardly.
+However, training without long thoughts might lead the model to a sub-optimal solution, like learning shortcuts.
+When adopting evaluation metrics that are not significantly dependent on the golden references (i.e., GEA, GRF, CometKiwi and GRB), DRT-14B shows its superior performance.
+Note that although GRB is a reference-based metric, it does not assess the model translations simply based on how similar they are to the golden references.
+DRT vs.
+Commercial LLMs.
+To give a deeper understanding of our DRT models’ performance, we further compare DRT models with GPT-4o OpenAI (2024a) and o1-preview OpenAI (2024b).
+The experimental results and corresponding analyses are provided in Appendix E.
+Model Flu. Sem. Lit.
+Qwen2.5-14B-Instruct -0.353 -0.363 -0.442
+QwQ-32B-Preview -0.063 0.022 -0.007
+Qwen2.5-14B-SFT 0.103 0.108 0.087
+DRT-14B 0.313 0.233 0.362
+Table 4: Human evaluation results in terms of fluency, semantic accuracy and literary quality.
+3.4 Human Evaluation
+We conduct human evaluation to further evaluate the performance of DRT-14B and strong baselines (Qwen2.5-14B-Instruct, QwQ-32B-Preview and Qwen2.5-14B-SFT).
+We randomly select 200 samples from our test set, and employ three human evaluators with high levels of fluency in English and Chinese to assess the generated translations from three aspects: fluency (Flu.), semantic accuracy (Sem.) and literary quality (Lit.).
+Following the Best-Worst Scaling method Kiritchenko and Mohammad (2017), evaluators are asked to select the best and the worst generated translation on each aspect.
+The result scores are calculated based on the percentage of times each model is selected as best minus the times it is selected as worst.
+Thus, the final scores should range from -1 (worst) to 1 (best).
+As shown in Table 4, DRT-14B significantly outperforms these strong baselines, especially in the aspect of literary quality.
+These results demonstrate the superiority of our DRT models.
+The Fleiss’ Kappa scores Fleiss (1971) of Flu., Sem.
+and Lit.
+are 0.75, 0.69 and 0.85, respectively, indicating a good inter-agreement among evaluators.
+3.5 Inference Time Analysis
+During evaluating LLMs’ literature translation performance on our test set, we leverage vLLM to accelerate the model generation.
+A single NVIDIA A100 GPU (40G) is used to deploy each LLM.
+As shown in Figure 5, the average time costs of DRT models are significantly higher than LLMs (w/o CoT).
+This is because DRT models should first generate the long thought and then provide the final translation, thus needing more inference time ($\times$11.9~13.9).
+This also indicates that the O1-like LLMs may not be applicable to some scenarios with high real-time requirements.
+Figure 5: Time cost during inference on the testing set.
+Source Sentence Qwen2.5-14B-Instruct QwQ-32B-Preview Qwen2.5-14B-SFT DRT-14B
 And if their words had taken another turn…if he himself had been less fastidious about intruding on another man’s secrets…it was cruel to think how thin a film had shut out rescue from all this guilt and misery.
-{CJK}UTF8gbsn 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 {CJK}UTF8gbsn 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 {CJK}UTF8gbsn 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
+{CJK}UTF8gbsn 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 {CJK}UTF8gbsn 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 {CJK}UTF8gbsn 如果他们的言辞稍有不同,他不那么介意侵扰别人的隐私,想到那层薄纱如何隔绝了救援,真是残忍。 {CJK}UTF8gbsn 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
 This cold officer upon a monument, who dropped epithets unconcernedly down, would be finer as a dead man, he thought.
-{CJK}UTF8gbsn 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 {CJK}UTF8gbsn 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 {CJK}UTF8gbsn 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
+{CJK}UTF8gbsn 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 {CJK}UTF8gbsn 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 {CJK}UTF8gbsn 这位冷酷的官员站在纪念碑上,随意地抛出贬义词,他觉得,作为死人会更显尊贵 {CJK}UTF8gbsn 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
 I see on your cheek two tears which I know are hot as two sparks, and salt as two crystals of the sea.
-{CJK}UTF8gbsn 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 {CJK}UTF8gbsn 我见你面颊两颗泪,温热如火花咸涩如海晶 {CJK}UTF8gbsn 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
+{CJK}UTF8gbsn 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 {CJK}UTF8gbsn 我见你面颊两颗泪,温热如火花咸涩如海晶 {CJK}UTF8gbsn 我见你脸颊上两颗泪珠,热如火花,咸似海盐。 {CJK}UTF8gbsn 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
 The night wore out, and, as he stood upon the bridge listening to the water as it splashed the river-walls of the Island of Paris, where the picturesque confusion of houses and cathedral shone bright in the light of the moon, the day came coldly, looking like a dead face out of the sky.
-{CJK}UTF8gbsn 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 {CJK}UTF8gbsn 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 {CJK}UTF8gbsn 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
+{CJK}UTF8gbsn 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 {CJK}UTF8gbsn 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 {CJK}UTF8gbsn 夜幕渐渐消散,他站在桥上,聆听水花轻拍巴黎岛的河岸。月光下,岛上房屋与大教堂交织成一幅迷人的画卷,显得格外明亮。清晨悄然而至宛如从天际浮现的一张冷峻的脸庞。 {CJK}UTF8gbsn 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
 Table 5: Case Studies of literature translation. Green indicates good translations, while red indicates bad ones.
-3.2 Comparison Models
-We use vanilla Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as the comparison models.
-Besides, two O1-like LLMs, i.e., Marco-o1-7B and QwQ-32B-preview, are also conducted as baselines.
-3.3 Main Results
-Table 2 shows the experimental results.
-After instruction tuning on the long-thought MT training data, DRT-series LLMs significantly outperform the corresponding vanilla backbones.
-For example, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 4.04 GRF, 1.42 CometKiwi, 4.46 GRB, 8.52 BLEU and 3.41 CometScore.
-DRT-o1-8B outperforms Llama-3.1-8B-Instruct by 5.24 GRF, 0.71 CometKiwi, 7.50 GRB, 14.12 BLEU and 4.23 CometScore.
-Besides, DRT-o1-14B outperforms all baselines (including QwQ-32B-preview and Marco-o1-7B) in terms of all metrics, showing the effectiveness of long thought in MT.
 3.6 Case Study
-Table 3 provides some literature translation cases of Qwen2.5-14B-Instruct, QwQ-32B-Preview and DRT-o1-14B.
-With the help of long thought, the translations of DRT-o1-14B align more closely with the conventions of the Chinese language and exhibit a greater literary quality.
-Besides, in addition to DRT-o1-14B, some translation snippets of other LLMs can also show a great performance (marked in green).
-This indicates that vanilla LLMs might have the capability to translate literature, and long thought could further activate the capability.
-Figure 4: Time cost during inference on the testing set.
-3.5 Inference Time Analysis
-When evaluating LLMs’ literature translation performance on our test set, we leverage vLLM to accelerate the model generation.
-Specifically, we use a single NVIDIA A100 GPU (40G) to deploy and call each LLM, and calculate the corresponding generation time in the test set.
-As shown in Figure 4, the average time costs of DRT-o1 models are significantly higher than vanilla LLMs.
-This is because DRT-o1 models should first generate the long thought and then provide the final translation, thus needing more inference time.
-This also indicates that the O1-like LLMs may not be applicable to some scenarios with high real-time requirements.
+Table 5 provides some literature translation cases of Qwen2.5-14B-Instruct, QwQ-32B-Preview, Qwen2.5-14B-SFT and DRT-14B.
+With the help of long thought, the translations of DRT-14B align more closely with the conventions of the Chinese language and exhibit a greater literary quality.
+In addition to DRT-14B, some translation snippets of other LLMs can also show a great performance (marked in green).
+This indicates that vanilla LLMs might have the capability to translate literature, and long thought could further activate this capability.
 4 Related Work
-4.1 O1-like LLMs
+O1-like LLMs.
 Recently, O1-like LLMs have shown great performance in reasoning tasks, especially math and coding tasks.
@@
 (2024) propose journey learning, a training paradigm, to encourage LLMs to learn not just shortcuts, but the complete exploration process.
-With only 327 training samples, the journey learning enhances LLMs’ long-thought ability.
 Huang et al.
@@
 (2024) is proposed to deal with open-ended text generation.
-Marco-o1 shows the effectiveness of a MCTS-enhanced inference method, and achieves great performance on math problems.
-Besides, Marco-o1 studies the machine translation cases with long thought, and shows the potentiality of O1-like models in dealing with MT.
-4.2 Literature Translation
+More recently, DeepSeek-R1 Guo et al.
+(2025) and Kimi K1.5 Team et al.
+(2025) are proposed, and show their promising reasoning ability.
+Literature Translation.
 Different from translating standard MT corpora (e.g., news articles), translating literature books is more difficult since it often requires equivalence beyond the word level Thai et al.
@@
 (2022) introduce Par3 to benchmark LLMs’ literature translation capability from non-English languages to English.
-Wang et al.
-(2024b) propose DelTA, an online document-level MT agent to translate book-level long documents based on short-term memory, long-term memory, summaries and proper noun records.
 5 Conclusion
-In this paper, we introduce DRT-o1, an attempt to bring the success of long-thought reasoning to neural machine translation (MT).
+In this paper, we introduce DRT, an attempt to bring the success of long-thought reasoning to neural machine translation (MT).
 Specifically, we synthesize the machine translation long-thought samples by a designed multi-agent framework and GPT-4o reformulation.
@@
 To synthesize the long thought machine translation process for these sentences, a translator, an advisor and an evaluator collaborate to translate the source sentence iteratively.
-GPT-4o is further employed to enhance the readability and fluency of the thought process collected via the multi-agent framework.
-Based on the synthesized data, we train DRT-o1 models (using Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as backbones), and show their effectiveness in literature translation.
-Appendix A GPT-4o Evaluator
+Based on the synthesized data, we train DRT models.
+Extensive experiments on literature translation demonstrate the effectiveness of DRT models in terms of automatic evaluation.
+Case study and human evaluation further verify their superiority.
+Limitations
+While we show the effectiveness of long thought in MT, there are some limitations worth noting: (1) We focus on English-to-Chinese translation in this work, and future work could extend the data and the method to other translation directions.
+(2) There is still a lack of accurate automatic evaluation metrics for literary translation.
+Previous literature translation work typically relies on human evaluation Fonteyne et al.
+(2020); Karpinska and Iyyer (2023), and points out that BLEU and Comet might not be suitable for evaluating literature translation Karpinska and Iyyer (2023).
+This is because literary translations carry the responsibility of both semantic and critical interpretation, as they must address the challenge of achieving equivalence that often extends beyond the level of individual words Thai et al.
+(2022).
+Ethical Considerations
+We discuss the main ethical considerations of DRT models as follows: (1) Copyright.
+We mine literature sentences from 400 English books provided by the Project Gutenberg public-domain book repository [^5] https://www.gutenberg.org/, where the books are typically more than fifty years old and their copyrights have expired.
+The book data also has been extracted and released by Kryscinski et al.
+(2022).
+Therefore, we can construct DRT data based on these books, and further release our synthesized data.
+(2) Licenses.
+We will release our model checkpoints and synthesized data under CC-BY-NC-SA 4.0 license.
+Appendix A Prompt in Data Synthesis
+A.1 Prompts in Literature Book Mining
+SYSTEM PROMPT: You are assigned to translate an English literary work into Chinese.
+The text may include descriptions or expressions that embody English cultural nuances, which may not resonate with Chinese language habits.
+In such instances, a literal translation may not be appropriate; instead, these sentences should be paraphrased to convey their intended meaning effectively.
+USER PROMPT: The English sentence is provided as follows: <english sentence> {sentence} </english sentence> Please begin by assessing whether the English sentence contains any metaphors or similes.
+If there are none, respond with "no metaphors and no similes." If the English sentence does contain metaphors or similes, provide a literal translation of them, and then evaluate whether the literal translation is appropriate and easy for Chinese natives to understand.
+If it is suitable, format your response as follows (two lines): "your literal translation for metaphors/similes here (in Chinese)" "suitable" If it is unsuitable, please provide the reason for the unsuitability.
+Format your response as follows (three lines): "your literal translation for metaphors/similes here (in Chinese)" "unsuitable" "reason for unsuitability here (in Chinese)"
+A.2 Prompts in Multi-Agent Framework
+Translator Agent (Word-level translation)
+Given an English sentence, identify the important words (usually nouns, verbs, technical terms, and named entities that require special attention in translation) and translate them into Chinese.
+Output the translations in JSON format, for example: {"EnglishWord1": "ChineseTranslation", "EnglishWord2": "ChineseTranslation"} The Chinese translations can be a single translation or multiple options as deemed appropriate.
+Translator Agent (Preliminary translation)
+SYSTEM PROMPT: Given an English sentence and a JSON object containing potential translations of important keywords, produce a Chinese literal translation of the entire sentence.
+Please directly output the Chinese translation without any descriptions.
+USER PROMPT: <English Sentence> {sentence} </English Sentence> <Potential Keyword Translation> {keyword translation} </Potential Keyword Translation>
+Translator Agent (Refinement translation)
+In the refine loop, the translator agent receives the feedback of the previous translation, and then provides a new translation.
+The prompt is a multi-turn dialogue between the translator and advisor, where the system prompt is the same as the preliminary translation.
+Advisor Agent
+Please rate the Chinese translation of the following English text and provide your comments and suggestions.
+Evaluator Agent
+SYSTEM PROMPT: Please evaluate the following Chinese translation of an English text.
+Rate the translation on a scale of 0 to 100, where: - 10 points: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a Chinese reader.
+- 30 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a Chinese reader to fully grasp the intended message.
+- 50 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a Chinese reader.
+- 70 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a Chinese reader.
+- 90 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a Chinese reader.
+Please provide the reason first, followed by a score.
+Format your evaluation in the JSON structure below: {"reason": "reason for the score", "score": int}
+A.3 Prompts in Thought Reformulation
+A student is engaged in the task of translating an English sentence into Chinese.
+The English sentence is as follows: <English Sentence> {sentence} </English Sentence> This student constantly thinks about and optimizes his translation.
+The whole process is shown as follows: <Translation Process> {translation process} </Translation Process> Please polish the whole translation process into a long first-person self-reflection description (use the present tense).
+The self-reflection should begin with selecting the keywords from the English sentence, translating the keywords, and then attempt to translate the whole sentence, and then think about whether the translation is good or not, and iteratively make translation attempts.
+Finally, make a final translation decision.
+Output the self-reflection description directly without any additional descriptions or explanations.
+Each line in the self-reflection description can be regarded as a reasoning step to the translation.
+Appendix B Details of Human Annotation
+In Section 2.4, we employ human annotation to provide the quality comparison between two translations for a source sentence.
+Specifically, we employ three Chinese master students with high levels of fluency in both English and Chinese as our human annotators.
+For each sample, we give the source sentence and its two translation (without the scores provided by our evaluator agent) to all three annotators, and every annotator should provide one of the following judgments: (1) the first translation is better than the second one; (2) the second translation is better than the first one; (3) two translation are similar in quality.
+During annotation, we encourage the annotators to give differentiated judgments, i.e., judgment (1) or judgment (2).
+If three annotators give the same judgment for a sample, the judgment will be regarded as the final label.
+Otherwise, its label will be decided by a group meeting of all three annotators and a senior data scientist.
+Appendix C GPT-4o Evaluator
 For GRB and GRF, we prompt GPT-4o (2024-08-06 version) as the MT evaluator in the reference-based and reference-free manners, respectively.
-The corresponding prompts borrow from Kocmi and Federmann (2023), and make some adaptions to literature translation (c.f.
-Figure 5).
-Appendix B Complete Generation of DRT-o1
-Figure 6 shows an example of DRT-o1-14B.
-As we can see, the model learns the thought process in our data collection.
-DRT-o1-14B first performs the word-level translation, and then attempts the preliminary translation.
-Next, it iteratively refines its translation until it thinks the translation is good enough.
+The corresponding prompts borrow from Kocmi and Federmann (2023), and make some adaptions to literature translation.
+GRB Prompt:
+Score the following translation from English to Chinese with respect to the human reference on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance".
+English source: {src} Chinese human reference: {ref} Chinese translation: {hyp} Score:
+GRF Prompt:
+Score the following translation from English to Chinese on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance".
+English source: {src} Chinese translation: {hyp} Score:
+Model reference-free reference-based
+GEA GRF CometKiwi GRB BLEU CometScore
+Commercial LLMs
+GPT-4o 71.88 85.57 73.01 82.78 34.51 79.41
+o1-preview 78.01 87.11 73.70 83.86 30.65 80.12
+DRT
+DRT-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
+DRT-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
+DRT-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
+Table 6: Experimental results of comparing DRT with commercial LLMs. The bold and the underline denote the best and second-best performances, respectively.
+Appendix D Implementation Details.
+Automatic Evaluation.
+To calculate CometKiwi and CometScore, we leverage the official codes [^6] https://github.com/Unbabel/COMET and the official models [^7] https://huggingface.co/Unbabel/wmt22-cometkiwi-da and https://huggingface.co/Unbabel/wmt22-comet-da.
+To calculate the BLEU score, we use the sacrebleu toolkit [^8] https://github.com/mjpost/sacrebleu to calculate the corpus-level BLEU.
+Training Details.
+Llama-Factory Zheng et al.
+(2024) is used to instruct-tune LLMs.
+All LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 8 (8$\times$1) batch size.
+We use the DeepSpeed ZeRO-3 optimization Rasley et al.
+(2020).
+Following Qin et al.
+(2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
+Inference Details.
+When evaluating model performance on the test set, we use vLLM toolkit Kwon et al.
+(2023) to accelerate the model generation.
+We use the sampling decoding strategy with 0.1 temperature, and set the repetition penalty to 1.05.
+For DeepSeek-R1 series (DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-14B and DeepSeek-R1-Distill-Qwen-32B), we follow the instruction [^9] https://github.com/deepseek-ai/DeepSeek-R1 to enforce them to avoid blank thinking.
+All experimental results listed in this paper are the average of 3 runs.
+Appendix E Comparison with Commercial LLMs
+As shown in Table 6, DRT-14B achieves competitive results with o1-preview, showing its superiority.
+Additionally, we observe that o1-preview significantly outperforms GPT-4o in terms of GEA.
+This finding highlights the effectiveness of long thought in machine translation.
+When applied to appropriate translation contexts, long thought can further enhance the authenticity of translations.

### v4

@@
 Using Qwen2.5 and LLama-3.1 as the backbones, DRT models can learn the thought process during machine translation, and outperform vanilla LLMs as well as LLMs which are simply fine-tuning on the paired sentences without long thought, showing its effectiveness.
+[^1] The synthesized data and model checkpoints are released at https://github.com/krystalan/DRT.
 DRT: Deep Reasoning Translation via Long Chain-of-Thought
-Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT[Uncaptioned image]DRT-7B[Uncaptioned image]DRT-8B[Uncaptioned image]DRT-14B
+Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com
 1 Introduction
@@
 There are two key points in achieving this goal:
-i) A suitable translation scenario to employ long thought in MT: Not all scenarios require long chain-of-thought (CoT)111“long CoT” is equal to “long thought”, and we alternatively use these two terms in this paper.
+i) A suitable translation scenario to employ long thought in MT: Not all scenarios require long chain-of-thought (CoT)222“long CoT” is equal to “long thought”, and we alternatively use these two terms in this paper.
 during translation.
@@
 To emphasize the valid thought process, translations without score change will be removed.
-That is, if $s^{i}$ is equal to $s^{i-1}$ ($i=1,2,...,m$), we will discard $\langle t^{i},f^{i},s^{i}\rangle$ in $\mathcal{P}(s)$, resulting in:
+That is, if $s^{i}$ is equal to $s^{i-1}$ (i=1,2,,mi=1,2,...,m), we will discard $\langle t^{i},f^{i},s^{i}\rangle$ in $\mathcal{P}(s)$, resulting in:
 $\begin{split}\mathcal{P}^{{}^{\prime}}(s):s\Rightarrow\langle\mathcal{W}^{\text{src}},\mathcal{W}^{\text{tgt}}\rangle\Rightarrow\langle t^{0},f^{0},s^{0}\rangle\\ \Rightarrow\langle t^{r_{1}},f^{r_{1}},s^{r_{1}}\rangle\Rightarrow...\Rightarrow\langle t^{r_{n}},f^{r_{n}},s^{r_{n}}\rangle\end{split}$ (2)
@@
 DRT
-DRT-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
-DRT-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
-DRT-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
-Table 3: Experimental results on literature translation.
-The bold and the underline denote the best and second-best performances, respectively.
-$\dagger$” and“$\ddagger$” denote statistically significant better than the corresponding SFT LLMs (w/o CoT) with t-test p < 0.01 and 0.05, respectively.
+DRT-8B $_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
+DRT-7B $_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
+DRT-14B $_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
+[…]
 3 Experiments
@@
 And if their words had taken another turn…if he himself had been less fastidious about intruding on another man’s secrets…it was cruel to think how thin a film had shut out rescue from all this guilt and misery.
-{CJK}UTF8gbsn 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 {CJK}UTF8gbsn 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 {CJK}UTF8gbsn 如果他们的言辞稍有不同,他不那么介意侵扰别人的隐私,想到那层薄纱如何隔绝了救援,真是残忍。 {CJK}UTF8gbsn 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
+而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 如果他们的言辞稍有不同,他不那么介意侵扰别人的隐私,想到那层薄纱如何隔绝了救援,真是残忍。 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
 This cold officer upon a monument, who dropped epithets unconcernedly down, would be finer as a dead man, he thought.
-{CJK}UTF8gbsn 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 {CJK}UTF8gbsn 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 {CJK}UTF8gbsn 这位冷酷的官员站在纪念碑上,随意地抛出贬义词,他觉得,作为死人会更显尊贵 {CJK}UTF8gbsn 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
+他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 这位冷酷的官员站在纪念碑上,随意地抛出贬义词,他觉得,作为死人会更显尊贵 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
 I see on your cheek two tears which I know are hot as two sparks, and salt as two crystals of the sea.
-{CJK}UTF8gbsn 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 {CJK}UTF8gbsn 我见你面颊两颗泪,温热如火花咸涩如海晶 {CJK}UTF8gbsn 我见你脸颊上两颗泪珠,热如火花,咸似海盐。 {CJK}UTF8gbsn 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
+我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 我见你面颊两颗泪,温热如火花咸涩如海晶 我见你脸颊上两颗泪珠,热如火花,咸似海盐。 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
 The night wore out, and, as he stood upon the bridge listening to the water as it splashed the river-walls of the Island of Paris, where the picturesque confusion of houses and cathedral shone bright in the light of the moon, the day came coldly, looking like a dead face out of the sky.
-{CJK}UTF8gbsn 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 {CJK}UTF8gbsn 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 {CJK}UTF8gbsn 夜幕渐渐消散,他站在桥上,聆听水花轻拍巴黎岛的河岸。月光下,岛上房屋与大教堂交织成一幅迷人的画卷,显得格外明亮。清晨悄然而至宛如从天际浮现的一张冷峻的脸庞。 {CJK}UTF8gbsn 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
+夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 夜幕渐渐消散,他站在桥上,聆听水花轻拍巴黎岛的河岸。月光下,岛上房屋与大教堂交织成一幅迷人的画卷,显得格外明亮。清晨悄然而至宛如从天际浮现的一张冷峻的脸庞。 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
 Table 5: Case Studies of literature translation. Green indicates good translations, while red indicates bad ones.
@@
 In such instances, a literal translation may not be appropriate; instead, these sentences should be paraphrased to convey their intended meaning effectively.
-USER PROMPT: The English sentence is provided as follows: <english sentence> {sentence} </english sentence> Please begin by assessing whether the English sentence contains any metaphors or similes.
-If there are none, respond with "no metaphors and no similes." If the English sentence does contain metaphors or similes, provide a literal translation of them, and then evaluate whether the literal translation is appropriate and easy for Chinese natives to understand.
-If it is suitable, format your response as follows (two lines): "your literal translation for metaphors/similes here (in Chinese)" "suitable" If it is unsuitable, please provide the reason for the unsuitability.
-Format your response as follows (three lines): "your literal translation for metaphors/similes here (in Chinese)" "unsuitable" "reason for unsuitability here (in Chinese)"
+USER PROMPT: The English sentence is provided as follows: ¡english sentence¿ {sentence} ¡/english sentence¿ Please begin by assessing whether the English sentence contains any metaphors or similes.
+If there are none, respond with ”no metaphors and no similes.” If the English sentence does contain metaphors or similes, provide a literal translation of them, and then evaluate whether the literal translation is appropriate and easy for Chinese natives to understand.
+If it is suitable, format your response as follows (two lines): ”your literal translation for metaphors/similes here (in Chinese)” ”suitable” If it is unsuitable, please provide the reason for the unsuitability.
+Format your response as follows (three lines): ”your literal translation for metaphors/similes here (in Chinese)” ”unsuitable” ”reason for unsuitability here (in Chinese)”
 A.2 Prompts in Multi-Agent Framework
@@
 Given an English sentence, identify the important words (usually nouns, verbs, technical terms, and named entities that require special attention in translation) and translate them into Chinese.
-Output the translations in JSON format, for example: {"EnglishWord1": "ChineseTranslation", "EnglishWord2": "ChineseTranslation"} The Chinese translations can be a single translation or multiple options as deemed appropriate.
+Output the translations in JSON format, for example: {”EnglishWord1”: ”ChineseTranslation”, ”EnglishWord2”: ”ChineseTranslation”} The Chinese translations can be a single translation or multiple options as deemed appropriate.
 Translator Agent (Preliminary translation)
@@
 Please directly output the Chinese translation without any descriptions.
-USER PROMPT: <English Sentence> {sentence} </English Sentence> <Potential Keyword Translation> {keyword translation} </Potential Keyword Translation>
+USER PROMPT: ¡English Sentence¿ {sentence} ¡/English Sentence¿ ¡Potential Keyword Translation¿ {keyword translation} ¡/Potential Keyword Translation¿
 Translator Agent (Refinement translation)
@@
 Please provide the reason first, followed by a score.
-Format your evaluation in the JSON structure below: {"reason": "reason for the score", "score": int}
+Format your evaluation in the JSON structure below: {”reason”: ”reason for the score”, ”score”: int}
 A.3 Prompts in Thought Reformulation
 A student is engaged in the task of translating an English sentence into Chinese.
-The English sentence is as follows: <English Sentence> {sentence} </English Sentence> This student constantly thinks about and optimizes his translation.
-The whole process is shown as follows: <Translation Process> {translation process} </Translation Process> Please polish the whole translation process into a long first-person self-reflection description (use the present tense).
+The English sentence is as follows: ¡English Sentence¿ {sentence} ¡/English Sentence¿ This student constantly thinks about and optimizes his translation.
+The whole process is shown as follows: ¡Translation Process¿ {translation process} ¡/Translation Process¿ Please polish the whole translation process into a long first-person self-reflection description (use the present tense).
 The self-reflection should begin with selecting the keywords from the English sentence, translating the keywords, and then attempt to translate the whole sentence, and then think about whether the translation is good or not, and iteratively make translation attempts.
@@
 GRB Prompt:
-Score the following translation from English to Chinese with respect to the human reference on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance".
+Score the following translation from English to Chinese with respect to the human reference on a continuous scale from 0 to 100, where score of zero means ”no meaning preserved” and score of one hundred means ”perfect preservation of meaning, with faithfulness, expressiveness, and elegance”.
 English source: {src} Chinese human reference: {ref} Chinese translation: {hyp} Score:
 GRF Prompt:
-Score the following translation from English to Chinese on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance".
+Score the following translation from English to Chinese on a continuous scale from 0 to 100, where score of zero means ”no meaning preserved” and score of one hundred means ”perfect preservation of meaning, with faithfulness, expressiveness, and elegance”.
 English source: {src} Chinese translation: {hyp} Score:
@@
 DRT
-DRT-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
-DRT-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
-DRT-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
+DRT-8B $_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
+DRT-7B $_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
+DRT-14B $_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
 Table 6: Experimental results of comparing DRT with commercial LLMs. The bold and the underline denote the best and second-best performances, respectively.
@@
 To calculate the BLEU score, we use the sacrebleu toolkit [^9] https://github.com/mjpost/sacrebleu to calculate the corpus-level BLEU.
-Training Details.
-Llama-Factory Zheng et al.
-(2024) is used to instruct-tune LLMs.
-All LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 8 (8$\times$1) batch size.
-We use the DeepSpeed ZeRO-3 optimization Rasley et al.
-(2020).
-Following Qin et al.
-(2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
+[…]
 Inference Details.